# SASSL: Enhancing Self-Supervised Learning via Neural Style Transfer

### 02 Dec 2023 v1

@@ 1 Introduction
Given the lack of labels, SSL relies on pretext tasks, i.e., predefined tasks where pseudo-labels can be generated. Some examples include contrastive learning Chen et al. (2020a); He et al. (2020), clustering Caron et al. (2021; 2020); Assran et al. (2022), and generative modeling He et al. (2022); Devlin et al. (2018). Many of these pretext tasks involve training the model to distinguish between different views of the same input and inputs corresponding to different samples. For these tasks, the way input data is augmented is crucial for the network to learn useful invariances and extract robust representations (Chen et al., 2020a). While state-of-the-art augmentations incorporate a wide range of primitive color, spectral and spatial transformations, they often disregard the natural structure of an image. As a result, SSL pre-training methods may generate augmented samples with degraded semantic information, and may be less able to capture diverse visual attributes.
@@ 2.1 Data Augmentation in SSL
Novel augmentation techniques have been proposed for both supervised and self-supervised learning. For example, Han et al. (2022) propose a novel technique to apply multiple data augmentation techniques in a local fashion by splitting the image into patches. This is aligned with other local augmentation techniques such as random erasing (Zhong et al., 2020), CutMix (Yun et al., 2019), Mixup (Zhang et al., 2017), among others. While these methods focus on masking or severely distorting image patches, disregarding their semantic information, our approach is a global transformation that decouples an image into its appearance and semantic attributes. This allows applying transformations exclusively to its appearance without affecting the image semantics.
@@ 3.2 Neural Style Transfer
Here, we assume ${\bm{z}}_{s}$ to be a contracted embedding of the style image ($D\ll H_{s}W_{s}$). The stylization network $\mathcal{T}$ is comprised by $L$ blocks $\{t_{l}\}_{l=1}^{L}$. These extract high-level features from the content image, align them to the style embedding ${\bm{z}}_{s}$ and map the resulting features back to the pixel domain.
@@ 4 Proposed Method: SASSL
SASSL operates over minibatches, allowing efficient data preprocessing. Let ${\bm{I}}_{c}\in\mathbb{R}^{B\times C\times H_{c}\times W_{c}}$ and ${\bm{I}}_{s}\in\mathbb{R}^{B\times C\times H_{s}\times W_{s}}$ be content and style minibatches, respectively, comprised by $B$ images ${\bm{I}}_{c}^{(b)}$ and ${\bm{I}}_{s}^{(b)},b\in\{1,\dots,B\}$. Then, the stylized minibatch ${\bm{I}}_{\text{cs}}\in\mathbb{R}^{B\times C\times H_{c}\times W_{c}}$ is generated by applying style transfer between a sample from the content batch and a sample from a style batch, given an arbitrary selection criterion. In what follows, we propose two alternatives for selecting style images to balance between augmentation diversity and efficiency.
@@ 5.1 Downstream Task Performance
Please refer to Section A.10 for additional ImageNet results showing SASSL’s performance on larger models (ResNet-50 4x pre-trained via MoCo v2) and different architectures (ViT-B$/16$ pre-trained via MoCo v3). We experimentally show that SASSL also improves downstream task performance under these scenarios, highlighting its potential across different model complexities and architectures.
@@
None (Baseline) $72.55\pm 0.67$ $53.23\pm 0.45$ $41.33\pm 0.06$ $\bf 75.92\pm 0.12$ $72.68\pm 0.7$
ImageNet (Ours) $74.07\pm 0.46$ $56.87\pm 0.43$ $45.01\pm 0.37$ $75.73\pm 0.12$ $73.69\pm 1.22$
@@
Retinopathy (Ours) $74.07\pm 0.62$ $\bf 56.99\pm 0.26$ $44.9\pm 0.16$ $75.78\pm 0.08$ $73.73\pm 0.57$
DTD (Ours) $74.32\pm 0.32$ $56.77\pm 0.36$ $\bf 45.08\pm 0.31$ $75.76\pm 0.11$ $\bf 74.41\pm 1.39$
PBN (Ours) $\bf 74.64\pm 0.43$ $56.76\pm 0.23$ $44.70\pm 0.37$ $75.75\pm 0.17$ $72.75\pm 1.01$
@@
None (Baseline) $74.89\pm 0.67$ $51.61\pm 0.13$ $77.92\pm 0.14$ $78.87\pm 0.2$ $71.54\pm 0.43$
ImageNet (Ours) $75.52\pm 0.28$ $51.74\pm 0.14$ $79.21\pm 0.07$ $79.64\pm 0.16$ $\bf 72.31\pm 1.87$
@@ 5.2 Transfer Learning Performance
Baseline -- -- 72.55±.67plus-or-minus72.55.6772.$5$67 $91.19\pm 0.34$
SASSL (Ours) Probability: $p=0.8,$ Blending: $\alpha\in[0.1,0.3]$ Interpolation $\beta=1$ ImageNet (in-batch) $70.87\pm 0.72$ $89.33\pm 0.34$
@@ 5.3 Ablation Study
To better understand the impact of each Style Transfer component in our augmentation pipeline, we conduct an ablation study over the hyperparameters of our style transfer model. Similar to previous experiments, we quantify our uncertainty by reporting the mean and variance obtained by five independent trials.
@@ Appendix A Appendix
  • Sec. A.7 includes ablation studies to better understand the effect of each component of NST in the downstream performance.
@@
  • Sec. A.9 covers recent work on semantic-aware data augmentation and neural style transfer, and discusses their differences with our proposed method.
  • Sec. A.10 shows additional experiments on downstream performance using a more complex representation model (ResNet-50 4x) and an alternative representation architecture (ViT-B$/16$).
@@ A.8 Computational Requirements
SASSL v2 $+$ MoCo v2 (Ours) $29.48$ $21.28$
@@ A.9 Comparison to Recent Data Augmentation Work
Zheng et al. (2019) utilize NST to create additional samples for supervised learning, using only eight style images. It’s important to note that their approach does not incorporate NST as data augmentation; instead, the stylized images are fixed and added to the original training dataset. Additionally, their experiments are limited to a single classification dataset (Caltech). This contrasts sharply with our method, which proposes different stylization strategies (external and in-batch) and ways to control the stylization effect (via ablation and interpolation) to augment data on-the-fly in order to enhance SSL performance. While both methods share the use of style NST, their tasks and approaches are distinct.
@@
Table 10: MoCo v2 + SASSL Downstream Performance using a ResNet-50 4x backbone. Top-1 and Top-5 classification accuracy obtained by pre-training and linear probing on ImageNet. Performance reported on a single random seed.
@@ A.10 Downstream Task Performance on Alternative Backbone Architectures
We ran additional experiments to explore the potential of SASSL on models with larger complexity and different architectures. In terms of complexity, we explored the use of ResNet-50 (4x) as representation model, which keeps the same architecture as the ResNet-50 models evaluated in Section 5, but increases the parameters from $23$ million to $375$ million. In terms of architecture, we evaluated the use of ViT-B$/16$, which is comprised by $86$ million parameters.
For comparison purposes, We pre-train and linear probe ResNet-50 (4x) representation models on ImageNet via MoCo v2 with its default augmentation and with SASSL. Both pre-training and downstream settings follow our default configuration, as documented in the Appendix Section A.3 and A.4. Similarly, we pre-train and linear probe ViT-B$/16$ representation models on ImageNet via MoCo v3 with its default augmentation and with SASSL. For the latter case, we updated the blending factor $\alpha$ to be uniformly sampled between $0.1$ and $0.5$ (instead of the default $0.1$ to $0.3$ interval). Note that SASSL does not introduce any additional training parameters.
Tables 10 and 11 report the downstream classification accuracy for ResNet-50 (x4) and ViT-B$/16$, respectively. ResNet-50 (4x) results show SASSL improves top-1 classification accuracy by $1.1\%$. This is consistent with our previous experiments, where ResNet-50 representation models pre-trained with SASSL improved accuracy by $2.1\%$. While still showing a significant improvement, we hypothesize that SASSL’s improvement drop from $2.1\%$ to $1.1\%$ is mainly because the large ResNet-50 (4x)’s model complexity. More precisely, ResNet-50 (4x) is capable of learning better representations than the original ResNet-50, so data augmentation strategies have less effect on its downstream classification accuracy.

### 02 Jan 2024 v2

@@ 1 Introduction
Given the lack of labels, SSL relies on pretext tasks, i.e., predefined tasks where pseudo-labels can be generated. Some examples include contrastive learning (Chen et al., 2020a; He et al., 2020), clustering (Caron et al., 2021; 2020; Assran et al., 2022), and generative modeling (He et al., 2022; Devlin et al., 2018). Many of these pretext tasks involve training the model to distinguish between different views of the same input and inputs corresponding to different samples. For these tasks, the way input data is augmented is crucial for the network to learn useful invariances and extract robust representations (Chen et al., 2020a). While state-of-the-art augmentations incorporate a wide range of primitive color, spectral and spatial transformations, they often disregard the natural structure of an image. As a result, SSL pre-training methods may generate augmented samples with degraded semantic information, and may be less able to capture diverse visual attributes.
@@ 2.1 Data Augmentation in SSL
Novel augmentation techniques have been proposed for both supervised and self-supervised learning. For example, Han et al. (2022) propose a novel method to apply multiple data augmentation techniques in a local fashion by splitting the image into patches. This is aligned with other local augmentation techniques such as random erasing (Zhong et al., 2020), CutMix (Yun et al., 2019), Mixup (Zhang et al., 2017), among others. While these methods focus on masking or severely distorting image patches, disregarding their semantic information, our approach is a global transformation that decouples an image into its appearance and semantic attributes. This allows applying transformations exclusively to its appearance without affecting the image semantics.
@@ 3.2 Neural Style Transfer
Here, we assume ${\bm{z}}_{s}$ to be a contracted embedding of the style image ($D\ll CH_{s}W_{s}$). The stylization network $\mathcal{T}$ is comprised by $L$ blocks $\{t_{l}\}_{l=1}^{L}$. These extract high-level features from the content image, align them to the style embedding ${\bm{z}}_{s}$ and map the resulting features to the pixel domain.
@@ 4 Proposed Method: SASSL
SASSL operates over minibatches, allowing efficient data preprocessing. Let ${\bm{I}}_{c}\in\mathbb{R}^{B\times C\times H_{c}\times W_{c}}$ and ${\bm{I}}_{s}\in\mathbb{R}^{B\times C\times H_{s}\times W_{s}}$ be content and style minibatches, respectively, comprised by $B$ images ${\bm{I}}_{c}^{(b)}$ and ${\bm{I}}_{s}^{(b)},b\in\{1,\dots,B\}$. Then, the stylized minibatch ${\bm{I}}_{\text{cs}}\in\mathbb{R}^{B\times C\times H_{c}\times W_{c}}$ is generated by applying style transfer between a sample from the content batch and a sample from the style batch, given an arbitrary selection criterion. In what follows, we propose two alternatives for selecting style images to balance between augmentation diversity and efficiency.
@@ 5.1 Downstream Task Performance
Please refer to Section A.10 for additional ImageNet results showing SASSL’s performance on larger models (ResNet-50 x4 pre-trained via MoCo v2) and different architectures (ViT-B$/16$ pre-trained via MoCo v3). We experimentally show that SASSL also improves downstream task performance under these scenarios, highlighting its potential across different model complexities and architectures.
@@
None (Baseline) $72.55\pm 0.67$ $53.23\pm 0.45$ $41.33\pm 0.2$ $\bf 75.88\pm 0.12$ $72.68\pm 0.7$
ImageNet (Ours) $74.07\pm 0.46$ $56.87\pm 0.43$ $45.01\pm 0.04$ $75.75\pm 0.1$ $73.69\pm 1.22$
@@
Retinopathy (Ours) $74.02\pm 0.61$ $\bf 56.99\pm 0.26$ $44.9\pm 0.16$ $75.78\pm 0.08$ $73.73\pm 0.57$
DTD (Ours) $74.32\pm 0.37$ $56.77\pm 0.36$ $\bf 45.08\pm 0.31$ $75.76\pm 0.11$ $\bf 74.41\pm 1.39$
PBN (Ours) $\bf 74.64\pm 0.43$ $56.9\pm 0.18$ $45.02\pm 0.14$ $75.79\pm 0.07$ $72.77\pm 0.77$
@@
None (Baseline) $74.89\pm 0.67$ $51.61\pm 0.13$ $77.92\pm 0.14$ $78.89\pm 0.2$ $71.54\pm 0.43$
ImageNet (Ours) $75.52\pm 0.23$ $51.74\pm 0.14$ $79.21\pm 0.07$ $79.64\pm 0.16$ $\bf 72.31\pm 1.85$
@@ 5.2 Transfer Learning Performance
MoCo v2 (default) -- -- 72.55±.67plus-or-minus72.55.6772.$5$67 $91.19\pm 0.34$
SASSL $+$ MoCo v2 (Ours) Probability: $p=0.8,$ Blending: $\alpha\in[0.1,0.3]$ Interpolation $\beta=1$ ImageNet (in-batch) $70.87\pm 0.72$ $89.33\pm 0.34$
@@ 5.3 Ablation Study
To better understand the impact of each proposed component, we conduct an ablation study over SASSL’s hyperparameters. Similar to previous experiments, we quantify our uncertainty by reporting the mean and variance obtained by five independent trials.
@@ Appendix A Appendix
  • Sec. A.7 includes ablation studies to better understand the effect of each component of Neural Style Transfer (NST) in the downstream performance.
@@
  • Sec. A.9 covers recent work on semantic-aware data augmentation and NST, and discusses their differences with our proposed method.
  • Sec. A.10 shows additional experiments on downstream performance using a more complex representation model (ResNet-50 x4) and an alternative representation architecture (ViT-B$/16$).
@@ A.8 Computational Requirements
SASSL $+$ MoCo v2 (Ours) $29.48$ $21.28$
@@ A.9 Comparison to Recent Data Augmentation Work
Zheng et al. (2019) utilize NST to create additional samples for supervised learning, using only eight style images. It’s important to note that their approach does not incorporate NST as data augmentation; instead, the stylized images are fixed and added to the original training dataset. Additionally, their experiments are limited to a single classification dataset (Caltech). This contrasts sharply with our method, which proposes different stylization strategies (external and in-batch) and ways to control the stylization effect (via ablation and interpolation) to augment data on-the-fly in order to enhance SSL performance. While both methods share the use of NST, their tasks and approaches are distinct.
@@
Table 10: MoCo v2 + SASSL Downstream Performance using a ResNet-50 x4 backbone. Top-1 and Top-5 classification accuracy obtained by pre-training and linear probing on ImageNet. Performance reported on a single random seed.
@@ A.10 Downstream Task Performance on Alternative Backbone Architectures
We ran additional experiments to explore the potential of SASSL on models with larger complexity and different architectures. In terms of complexity, we explored the use of ResNet-50 (x4) as representation model, which keeps the same architecture as the ResNet-50 models evaluated in Section 5, but increases the parameters from $23$ million to $375$ million. In terms of architecture, we evaluated the use of ViT-B$/16$, which is comprised by $86$ million parameters.
For comparison purposes, We pre-train and linear probe ResNet-50 (x4) representation models on ImageNet via MoCo v2 with its default augmentation and with SASSL. Both pre-training and downstream settings follow our default configuration, as documented in the Appendix Section A.3 and A.4. Similarly, we pre-train and linear probe ViT-B$/16$ representation models on ImageNet via MoCo v3 with its default augmentation and with SASSL. For the latter case, we updated the blending factor $\alpha$ to be uniformly sampled between $0.1$ and $0.5$ (instead of the default $0.1$ to $0.3$ interval). Note that SASSL does not introduce any additional training parameters.
Tables 10 and 11 report the downstream classification accuracy for ResNet-50 (x4) and ViT-B$/16$, respectively. ResNet-50 (x4) results show SASSL improves top-1 classification accuracy by $1.1\%$. This is consistent with our previous experiments, where ResNet-50 representation models pre-trained with SASSL improved accuracy by $2.1\%$. While still showing a significant improvement, we hypothesize that SASSL’s improvement drop from $2.1\%$ to $1.1\%$ is mainly because the large ResNet-50 (x4)’s model complexity. More precisely, ResNet-50 (x4) is capable of learning better representations than the original ResNet-50, so data augmentation strategies have less effect on its downstream classification accuracy.